OpenStudy (anonymous):

Please someone can check if my answers #2, #4, #6 and #8 are correct and some help with #10. Thanks

OpenStudy (anonymous):

Could you post the answers and the question so we can help? :)

OpenStudy (anonymous):

Are you not sure why you got them wrong? :-)

OpenStudy (anonymous):

*if

OpenStudy (anonymous):

So why don't we work each one of these out and then check if you got them right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Alright so #2 is \[\frac{ 7 }{ 10 }-\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

So we need to find a common denominator to subtract. Do you know what that would be? :)

OpenStudy (anonymous):

20

OpenStudy (anonymous):

Well, a simpler way would be to just make 1/2= 5/10. So you have 7/10-5/10=2/10, or 1/5 :) But if you make the common denominator 20, that would make it be 14/20 - 10/20= 4/20, or 1/5 :)

OpenStudy (anonymous):

So #2 is correct :o)

OpenStudy (anonymous):

#4 the common denominator is 40

OpenStudy (anonymous):

Yep, #4 is correct :-)

OpenStudy (anonymous):

#6 is right as well

OpenStudy (anonymous):

And so is #8! Good job :)

OpenStudy (anonymous):

I need help with # 10

OpenStudy (anonymous):

Ok let me look at it real quick, what do you need help understanding? :)

OpenStudy (anonymous):

For #10, find a common denominator (hint, multiply 7/8 by 10, and 9/10 by 8) and add the two together to get your total :)

OpenStudy (anonymous):

the common denominator is 80. 7x10=70 9x8=72 . 70+72=142. now we divide 142/80= 1.62/80

OpenStudy (anonymous):

so the answer is 1 cup 62/80

OpenStudy (anonymous):

Yep! Now you can reduce that :)

OpenStudy (anonymous):

reduce what ?

OpenStudy (anonymous):

Your answer, 62/80. Divide 62 by 2 and 80 by 2. What do you get?

OpenStudy (anonymous):

ok so the answer is 1 cup 32/40

OpenStudy (anonymous):

Yep, did you know you can reduce that by 8? 32/8= 4 40/8=5 So the asnwer is 4/5 :)

OpenStudy (anonymous):

*answer

OpenStudy (anonymous):

1cup 4/5

OpenStudy (anonymous):

Yep! Nice job :)

OpenStudy (anonymous):

Thanks Kit_Kat_Nat_<3 for your help

OpenStudy (anonymous):

You are welcome, blanca, I'm happy to help! ^-^